OpenStudy (anonymous):

how to solve this equation? 4^x=3^(x+10)?

OpenStudy (loser66):

log both sides first use property of log to take the exponent down solve as usual

OpenStudy (anonymous):

I did. log4^x=log3^(x+10) then xlog4=(x+10)log3. Can you tell me how to have x on both sides?

OpenStudy (loser66):

the right hand side is = x log 3 +10 log 3 , right?

OpenStudy (anonymous):

Sorry how to put x on only one side without any other junk.

OpenStudy (loser66):

got my recent post?

OpenStudy (anonymous):

yes

OpenStudy (loser66):